# reference.bib

### v1

@@ Abstract.
We investigate the analytic properties of a Dirichlet series involving the Fourier-Jacobi coefficients of two cusp forms for orthogonal groups of signature $(2,n+2)$. Using an orthogonal Eisenstein series of Klingen type, we obtain an integral representation for this Dirichlet series. In the case when the corresponding lattice has only one 11 11-dimensional cusp, we rewrite this Eisenstein series in the form of an Epstein zeta function. If additionally $4\mid n$, we deduce a theta correspondence between this Eisenstein series and a Siegel Eisenstein series for the symplectic group of degree 22 22. We obtain, in this way, the meromorphic continuation of the Dirichlet series to $\mathbb{C}$ as a corollary. In the case of the $E_{8}$ lattice, we are able to further deduce a precise functional equation for the Dirichlet series.
@@ Acknowledgements.
The author was supported by the HIMR/UKRI "Additional Funding Programme for Mathematical Sciences", grant number EP/V521917/1 as well as by a scholarship by Onassis Foundation. The author would also like to thank Prof. Thanasis Bouganis for suggesting this problem as well as for his continuous guidance and support.
@@ 1. Introduction
$\mathcal{D}_{F,G}(s)=\zeta(2s-2k+4)\sum_{N=1}^{\infty}\langle\phi_{N},\psi_{N}\rangle N^{-s},$
[…]
@@
We wish to analyse their first result in greater detail. The method for their proof can be summarized in two steps: The first (and easier one) is to obtain an integral representation for $\mathcal{D}_{F,G}$, using a non-holomorphic Eisenstein series of Klingen type. The second is then to prove the meromorphic continuation and functional equation for this Eisenstein series. The proof of that involves writing the Eisenstein series in a form of an Epstein zeta function and then proving that it is a Mellin transform of a specific theta series.
@@
This general method of proof has been successful in a number of other cases as well. For example, Raghavan and Sengupta in [raghavan] and Gritsenko in [gritsenko] considered the same problem, but in the case when $F,G$ are Hermitian cusp forms of degree 22 22 over $\mathbb{Q}(i)$. In both papers, the authors managed to deduce the analytic properties of $\mathcal{D}_{F,G}$ by applying a very similar idea, however with two key differences regarding the second step above. The first one is that the Klingen Eisenstein series arises as the inner product of a theta series and a classical Eisenstein series for $\hbox{SL}_{2}$. The second and more important one is that it is now necessary to apply some differential operators on the theta series first. The reason for this is that there are terms that cause the inner-product integral to diverge, so we need to eliminate them with the use of differential operators.
@@
It should be noted here that the degree of the modular objects considered is not important. Yamazaki in [yamazaki] generalised the result of Kohnen and Skoruppa for Siegel cusp forms of arbitrary degree $n$. However, his method of deducing the analytic properties of the Eisenstein series of Klingen type is derived from the general Langlands’ theory. Krieg on the other hand in [krieg1991], used theta correspondence in the arbitrary degree $n$ case for Siegel, Hermitian (over $\mathbb{Q}(i)$) and quaternionic (over the Hamiltonian quaternions) cusp forms. The use of differential operators was again essential for the last two cases. What is even more interesting, is that Deitmar and Krieg in [deitmar_krieg], managed to prove theta correspondence of Eisenstein series for the groups $\hbox{Sp}_{n}(\mathbb{Z})$ and $\hbox{O}(m,m)$, for arbitrary $m,n\geq 1$. Their proof is based on proving the existence of an invariant differential operator $R$, which when applied to the suitable theta series, eliminates the terms that cause the divergence of the inner-product integral. This was a major advance, because up until then, all the operators had to be found explicitly and this could only be done in a handful of cases.
@@
In this paper, we consider a Dirichlet series of the same type, but for two cups forms $F,G$ of orthogonal groups of real signature $(2,n+2),\textup{ }n\geq 1$. Our purpose is to deduce the analytic continuation and functional equation of $\mathcal{D}_{F,G}$ in this case. The question of whether this series can be related to $L$-functions has been addressed in our previous paper [psyroukis_orthogonal]. Our motivation is the accidental isogenies of orthogonal groups of signature $(2,n+2)$ with classical groups for small $n$ (e.g. $\hbox{SO}(2,3)$ is isogenous to $\hbox{Sp}_{2}$ and $\hbox{SO}(2,4)$ is isogenous to $\hbox{U}(2,2)$). Therefore, already the results of Kohnen and Skoruppa in [kohnen_skoruppa] and Gritsenko in [gritsenko] should indicate a connection in the $n=1,2$ case. In principal, we use the same general method of proof as all the aforementioned papers (except Yamazaki in [yamazaki]). In our situation, we essentially obtain a theta-correspondence between Eisenstein series for $\hbox{Sp}_{2}$ and $\hbox{SO}(2,n+2)$.
@@
Let us now give a brief description of our results. The setup is as follows: We start with an even positive definite symmetric matrix $S$ of rank $n\geq 1$ so that the lattice $L:=\mathbb{Z}^{n}$ is maximal. We then set
$S_{0}=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\textup{ }S_{1}=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
@@
$G^{*}_{K}=\{g\in\hbox{SL}_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},\textup{ }G_{K}=\{g\in\hbox{SL}_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
@@
In order to apply the method we mentioned above, let for $Z\in\mathcal{H}_{S}$ and $\textup{Re}(s)>n+1$, $E(Z,s)$ denote an Eisenstein series of Klingen type in this setting (see Definition 4.1). The first step, namely the integral representation of the Dirichlet series, is then Proposition 4.2.
@@
For $Z\in\mathcal{H}_{S}$ and $s\in\mathbb{C}$ with $\textup{Re}(s)>n+2$, we have
$\langle F(Z)E(Z,s),G(Z)\rangle=\frac{1}{\#\hbox{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$
[…]
$\hbox{SO}(S;\mathbb{Z}):=\{g\in\hbox{SL}_{n}(\mathbb{Z})\mid g^{t}Sg=S\},$
which is a finite group.
For our next step, namely the theta-correspondence of Eisenstein series, we need several preparations. The first one is to rewrite the Eisenstein series in a form similar to an Epstein zeta function. We are able to do this only in the case when $\Gamma_{S}$ has one 11 11-dimensional cusp (see Definition 5.1). In this case, we have Proposition 5.1.
Proposition 5.1.
Let $S$ be such that #$\#\mathcal{C}^{1}(\Gamma_{S})=1$. Then, for each $Z\in\mathcal{H}_{S}$, there is a well-defined $R_{Z}$ in the space of majorants of $S_{1}$ (see Definition 5.2) such that
$E(Z,s)=\sum_{\ell\in X/\textup{GL}_{2}(\mathbb{Z})}\left(\det(R_{Z}[\ell])\right)^{-s/2},$
@@
Let now $Z\in\mathbb{H}_{2}$ (Siegel’s upper half plane, see section 6) and $W\in\mathcal{H}_{S}$. We define the theta series $\Theta(Z,W)$ as in Proposition 6.3. This is a real analytic function that transforms as a degree 22 22 Siegel modular form of weight $k=-n/2$ and a (specified) character $\chi$, under the action of the congruence subgroup $\Gamma_{0}^{(2)}(q)$ of $\hbox{Sp}_{2}(\mathbb{Z})$ (see Proposition 6.1). Here, $q$ denotes the level of $S$ (see Definition 2.4).
The next step is to consider the inner product of this theta series with a Siegel Eisenstein series for $\hbox{Sp}_{2}$ and hopefully obtain $E(Z,s)$. However, we cannot do this directly, because this theta series contains terms that would cause the inner-product integral against a hypothetical appropriate symplectic Eisenstein series to diverge. These are the summands that do not have full rank. It is therefore necessary to apply suitable differential operators in order to eliminate those terms.
It turns out that the $\hbox{Sp}_{2}(\mathbb{R})$-invariant differential operator $R$, defined in [deitmar_krieg], can be applied in this case too. However, $\Theta$ is not invariant under $\hbox{Sp}_{2}$ (or more accurately under its congruence subgroup $\Gamma_{0}^{(2)}(q)$) and therefore, applying $R$ to it would not be helpful. For this reason, we first apply the so-called Maass-Shimura operator $\delta_{k}^{(r)}$ (see Section 7). This, when applied to a modular function, increases its weight by $2r$. As $\Theta$ transforms as a weight $-n/2$ modular form, we have to pick $r=n/4$ and therefore it is necessary to impose the condition $4\mid n$. After taking all these factors into account, we obtain the main Theorem 8.1.
@@
Let $\widetilde{E}(Z,\chi,s)$ denote the symplectic Siegel Eisenstein series, as in Definition 8.1. Let $S$ have rank $n$, with $4\mid n$ and be such that #$\#\mathcal{C}^{1}(\Gamma_{S})=1$. Let also $k=-n/2$ and $r=n/4$, as before. We then have for $\hbox{Re}(s)>n+1$
@@
[…]
The meromorphic continuation of $\mathcal{D}_{F,G}$ to $\mathbb{C}$ follows then as a corollary (see Corollary 1).
As an application, when $S$ corresponds to the $E_{8}$ lattice, we are able to deduce a precise functional equation for the Dirichlet series (see Theorem 9.1). We choose this specific lattice because it satisfies the conditions of our main Theorem 8.1 and is the unique unimodular even lattice with only one 11 11-dimensional cusp. Hence, in this case, the symplectic Eisenstein series corresponds to the full $\hbox{Sp}_{2}(\mathbb{Z})$ and therefore a functional equation is available.
The structure of the paper is as follows: In sections 2 and 3, we give the definitions of the main objects of consideration, i.e. modular forms, Fourier-Jacobi forms and the Dirichlet series of interest. In section 4, we obtain the integral representation of the Dirichlet series and in section 5, we rewrite the Eisenstein series of Klingen type in a form of an Epstein zeta function. In section 6, we define the appropriate theta series for our case and section 7 is devoted in discussing the differential operators needed in order to obtain the theta-correspondence. This is done in section 8. Finally, in section 9, we obtain a functional equation in the case of the $E_{8}$ lattice.
@@
For a complex number $z$, we denote by $e(z):=e^{2\pi iz}$. For any vector $v$, we denote by $v^{t}$ its transpose. We also use the bracket notation $A[B]:=\overline{B}^{t}AB$ for suitably sized complex matrices $A,B$. Let $\zeta(s)$ denote the Riemann zeta function and $\Gamma(s):=\int_{0}^{\infty}t^{s-1}e^{-t}\hbox{d}t$ denote the gamma function.
@@ 2. Preliminaries
Let $V$ denote a finite dimensional vector space over $\mathbb{Q}$. We start with the following definition:
@@
A $\mathbb{Z}$-lattice $\Lambda$ is a free, finitely generated $\mathbb{Z}$-module, which spans $V$ over $\mathbb{Q}$, together with a symmetric non-degenerate bilinear form $V\times V\longrightarrow\mathbb{Q}$.
In the following, let $V:=\mathbb{Q}^{n}$ and $L:=\mathbb{Z}^{n}$ with $n\geq 1$. Assume $S$ is an even integral positive definite symmetric matrix of rank $n\geq 1$. Even here means $S[x]\in 2\mathbb{Z}$ for all $x\in L$. Write also $Q=S/2$.
Let then $V_{0}:=\mathbb{Q}^{n+2}$ and consider
$S_{0}=\begin{pmatrix}0&0&1\\ 0&-S&0\\ 1&0&0\end{pmatrix}$
of real signature $(1,n+1)$ and write $Q_{0}=S_{0}/2$. Define also the bilinear form
$\displaystyle\phi_{0}:V_{0}\times V_{0}$ $\displaystyle\longmapsto\mathbb{Q}$
$\displaystyle(x,y)$ $\displaystyle\longmapsto\frac{1}{2}x^{t}S_{0}y.$
Consider now a Witt decomposition of the form
$V_{0}=e_{n+1}\perp V\perp e_{1},$
where $e_{i}$ is the $i$’th standard basis vector. We then have that $\phi:=\phi_{0}\mid_{V\times V}$ is just
$(x,y)\longmapsto-\frac{1}{2}x^{t}Sy.$
Moreover, define $S_{1}=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$ and define the corresponding special orthogonal groups:
$G^{*}_{\mathbb{Q}}:=\{g\in\textup{SL}_{n+2}(\mathbb{Q})\mid g^{t}S_{0}g=S_{0}\},$
$G_{\mathbb{Q}}:=\{g\in\textup{SL}_{n+4}(\mathbb{Q})\mid g^{t}S_{1}g=S_{1}\}.$
We view $G^{*}_{\mathbb{Q}}$ as a subgroup of $G_{\mathbb{Q}}$ via the embedding $g\longmapsto\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}$.
We also use the notation $\hbox{SO}^{\phi_{0}}(V_{0})$ and in general $\hbox{SO}^{\phi}(V)$ for any quadratic space $(V,\phi)$.
@@
For a matrix $g\in\textup{Mat}(n+4,\mathbb{R})$, we write it as
@@
with $A\in\textup{Mat}(n+2,\mathbb{R}),\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $a,b,c,d$ real column vectors. Now the map
$Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$
@@
Let also $\Gamma^{+}(L_{0})=\Gamma(L_{0})\cap G^{*,0}_{\mathbb{R}}$. If then
@@
where $Z\in\mathcal{H}_{S}$. It is then Koecher’s principle that gives us that $A(r)=0$ unless $r\in L_{0}^{*}\cap\overline{\mathcal{P_{S}}}$ (cf. [eisenstein_thesis, Theorem 1.5.2]). Here, $\overline{\mathcal{P}_{S}}$ denotes the closure of $\mathcal{P_{S}}$. By [eisenstein_thesis, Theorem 1.6.23], we have the following definition for cusp forms.
@@
$\langle F,G\rangle:=\int_{\mathcal{Q}}F(Z)\overline{G(Z)}\left(Q_{0}[\textup{Im}Z]\right)^{k}\hbox{d}^{*}Z,$
where $\hbox{d}^{*}Z=\left(Q_{0}[\textup{Im}Z]\right)^{-(n+2)}\hbox{d}Z$ denotes the $G_{\mathbb{R}}^{0}-$invariant volume element on $\mathcal{H}_{S}$. We remind here again that $Q_{0}=S_{0}/2$.

### v2

# Analytic Properties of an orthogonal Fourier-Jacobi Dirichlet Series

@@ Abstract.
We investigate the analytic properties of a Dirichlet series involving the Fourier-Jacobi coefficients of two cusp forms for orthogonal groups of signature $(2,n+2)$. Using an orthogonal Eisenstein series of Klingen type, we obtain an integral representation for this Dirichlet series. In the case when the corresponding lattice has only one $1$-dimensional cusp, we rewrite this Eisenstein series in the form of an Epstein zeta function. If additionally $4\mid n$, we deduce a theta correspondence between this Eisenstein series and a Siegel Eisenstein series for the symplectic group of degree $2$. We obtain, in this way, the meromorphic continuation of the Dirichlet series to $\mathbb{C}$ as a corollary. In the case of the $E_{8}$ lattice, we are able to further deduce a precise functional equation for the Dirichlet series.
Keywords: orthogonal modular forms, Fourier-Jacobi forms, Dirichlet series, theta correspondence, differential operators.
@@ Acknowledgements.
This work was supported by the Additional Funding Programme for Mathematical Sciences, delivered by EPSRC (EP/V521917/1) and the Heilbronn Institute for Mathematical Research, as well as by a scholarship from the Onassis Foundation. The author would also like to thank Prof. Thanasis Bouganis for his continuous guidance and support.
2020 Mathematics Subject Classification:
11F03, 11F60, 11F66
@@ 1. Introduction
(1) $D_{F,G}(s)=\zeta(2s-2k+4)\sum_{N=1}^{\infty}\langle\phi_{N},\psi_{N}\rangle N^{-s},$
where $\phi_{N},\psi_{N}$ are the $N$th Fourier-Jacobi coefficients of two degree two Siegel cusp forms $F,G$ of integral weight $k\geq 0$ and $\langle\;,\;\rangle$ denotes the Petersson inner product on the space of Jacobi forms of index $N$ and weight $k$. This can be formally seen as the analogue of the Rankin-Selberg convolution series for elliptic modular forms. They prove two main results: Firstly, this series admits a meromorphic continuation to $\mathbb{C}$ and satisfies a functional equation. Secondly, when $F$ is a Hecke eigenform and $G$ is in the Maass space, this series is proportional to the spinor $L$-function attached to $F$.
@@
We wish to analyse their first result in greater detail. The method for their proof can be summarised in two steps: the first (and easier) is to obtain an integral representation for $D_{F,G}$ using a non-holomorphic Eisenstein series of Klingen type, and the second is to prove the meromorphic continuation and functional equation for this Eisenstein series. The proof of that involves writing the Eisenstein series in the form of an Epstein zeta function and then proving that it is a Mellin transform of a specific theta series.
@@
This general method of proof has been successful in a number of other cases as well. For example, Raghavan and Sengupta in [raghavan] and Gritsenko in [gritsenko] considered the same problem, but in the case when $F,G$ are Hermitian cusp forms of degree $2$ over $\mathbb{Q}(i)$. In both papers, the authors managed to deduce the analytic properties of (a slightly differently normalised) $D_{F,G}$ by applying a very similar idea; however, there are two key differences in the second step above. The first one is that the Klingen Eisenstein series arises as the inner product of a theta series and a classical Eisenstein series for $\textup{SL}_{2}$. The second and more important one is that it is now necessary to apply some differential operators to the theta series first. The reason for this is that there are terms that cause the inner-product integral to diverge, so we need to eliminate them with the use of differential operators.
@@
It should be noted here that the degree of the modular objects considered is not important. Yamazaki in [yamazaki] generalised the result of Kohnen and Skoruppa for Siegel cusp forms of arbitrary degree $n\geq 1$. However, his method for deducing the analytic properties of the Eisenstein series of Klingen type is derived from the general Langlands’ theory. Krieg, on the other hand, in [krieg1991], used theta correspondence in the arbitrary degree $n$ case for Siegel, Hermitian (over $\mathbb{Q}(i)$) and quaternionic (over the Hamiltonian quaternions) cusp forms. The use of differential operators was again essential for the last two cases. What is even more interesting is that Deitmar and Krieg in [deitmar_krieg] managed to prove theta correspondence of Eisenstein series for the groups $\textup{Sp}_{n}(\mathbb{Z})$ and $\textup{O}(m,m)$, for arbitrary $m,n\geq 1$. Their proof is based on proving the existence of an invariant differential operator $R$, which, when applied to the suitable theta series, eliminates the terms that cause the divergence of the inner-product integral. This was a major advance because, up until then, all operators had to be found explicitly, and this could be done in only a handful of cases.
@@
In this paper, we consider a non-normalised Dirichlet series $\mathcal{D}_{F,G}$ of the same type as (1) for two cusp forms $F,G$ of orthogonal groups of real signature $(2,n+2),\textup{ }n\geq 1$. Our purpose is to deduce its analytic continuation and functional equation in this case. The question of whether this series can be related to $L$-functions has been addressed in our previous paper [psyroukis_orthogonal]. Our motivation is the accidental isogenies of orthogonal groups of signature $(2,n+2)$ with classical groups for small $n$ (e.g. $\textup{SO}(2,3)$ is isogenous to $\textup{Sp}_{2}$ and $\textup{SO}(2,4)$ is isogenous to $\textup{U}(2,2)$). Therefore, the results of Kohnen and Skoruppa in [kohnen_skoruppa] and of Gritsenko in [gritsenko] already indicate a connection in the $n=1,2$ case. In principle, we use the same general method of proof as all the aforementioned papers (except Yamazaki in [yamazaki]). In our situation, we essentially obtain a theta-correspondence between Eisenstein series for $\textup{Sp}_{2}$ and $\textup{SO}(2,n+2)$.
@@
Let us now give a brief description of our results. The setup is as follows: We start with an even positive definite symmetric matrix $S$ of rank $n\geq 1$ so that the lattice $L:=\mathbb{Z}^{n}$ is $\mathbb{Z}$-maximal. We then set
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\textup{ }S_{1}:=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
@@
$G^{*}_{K}:=\{g\in\textup{SL}_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},\textup{ }G_{K}:=\{g\in\textup{SL}_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
@@
where $\langle\;,\;\rangle$ is a suitable inner product defined on the space of Fourier-Jacobi forms of certain weight and (lattice) index (see Definition 3.6). This series converges for $\textup{Re}(s)>k+1$ (see Lemma 4.1).
In order to apply the method we mentioned above, for $W\in\mathcal{H}_{S}$ and $\textup{Re}(s)>n+1$, let $E(W,s)$ denote an Eisenstein series of Klingen type in this setting (see Definition 4.2). The first step, namely the integral representation of the Dirichlet series, is then Proposition 4.4.
Proposition 4.4.
For $W\in\mathcal{H}_{S}$ and $s\in\mathbb{C}$ with $\textup{Re}(s)>n+2$, we have
$\langle F(W)E(W,s),G(W)\rangle=\frac{1}{\#\textup{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$
@@
$\textup{SO}(S;\mathbb{Z}):=\{g\in\textup{SL}_{n}(\mathbb{Z})\mid g^{t}Sg=S\},$
which is a finite group.
For our next step, namely the theta-correspondence of Eisenstein series, we need several preparations. The first one is to rewrite the Eisenstein series in a form similar to an Epstein zeta function. We are able to do this only in the case when $\Gamma_{S}$ has one $1$-dimensional cusp (see Definition 5.1). In this case, we have Proposition 5.3.
@@
Let $S$ be such that $\#\mathcal{C}^{1}(\Gamma_{S})=1$. Then, for each $W\in\mathcal{H}_{S}$, there is a well-defined $R_{W}$ in the space of majorants of $S_{1}$ (see Definition 5.2) such that
$E(W,s)=\sum_{\ell\in X/\textup{GL}_{2}(\mathbb{Z})}\left(\det(R_{W}[\ell])\right)^{-s/2},$
@@
Let now $Z\in\mathbb{H}_{2}$ (Siegel’s upper half plane, see Section 6) and $W\in\mathcal{H}_{S}$. We define the theta series $\Theta(Z,W)$ as in Proposition 6.3. This is a real analytic function that transforms as a degree $2$ Siegel modular form of weight $k=-n/2$ and a (specified) character $\chi$, under the action of the congruence subgroup $\Gamma_{0}^{(2)}(q)$ of $\textup{Sp}_{2}(\mathbb{Z})$ (see Proposition 6.5). Here, $q$ denotes the level of $S$ (see Definition 2.4).
The next step is to consider the inner product of this theta series with a Siegel Eisenstein series for $\textup{Sp}_{2}$ and hopefully obtain $E(W,s)$. However, we cannot do this directly because the theta series contains terms that would cause the inner product integral against a hypothetical appropriate symplectic Eisenstein series to diverge. These are the summands that do not have full rank. It is therefore necessary to apply suitable differential operators in order to eliminate those terms.
It turns out that the $\textup{Sp}_{2}(\mathbb{R})$-invariant differential operator $R$, defined in [deitmar_krieg], can be applied in this case too. However, $\Theta$ is not invariant under $\textup{Sp}_{2}$ (or more accurately under its congruence subgroup $\Gamma_{0}^{(2)}(q)$) and therefore, applying $R$ to it would not be helpful. For this reason, we first apply the so-called Maass-Shimura operator $\delta_{k}^{(r)}$ (see (12)). This, when applied to a modular function, increases its weight by $2r$. As $\Theta$ transforms as a weight $-n/2$ modular form, we have to pick $r=n/4$, and therefore it is necessary to impose the condition $4\mid n$. After taking all these factors into account, we obtain the main Theorem 8.2.
@@
Let $\widetilde{E}(Z,\chi,s)$ denote the symplectic Siegel Eisenstein series, as in Definition 8.1. Let $S$ have rank $n$, with $4\mid n$ and be such that $\#\mathcal{C}^{1}(\Gamma_{S})=1$. Let also $k=-n/2$ and $r=n/4$, as before. We then have for $\textup{Re}(s)>n+1$
@@
where $\gamma_{S}(s)$ is an explicit gamma factor, $\xi(s)$ is the completed zeta function and  , $\langle\textup{ },\textup{ }\rangle_{\Gamma_{0}^{(2)}(q)}$ denotes the symplectic inner product of Definition 6.2.
The meromorphic continuation of $\mathcal{D}_{F,G}$ to $\mathbb{C}$ follows then as a corollary (see Corollary 8.4).
As an application, when $S$ corresponds to the $E_{8}$ lattice, we are able to deduce a precise functional equation for the Dirichlet series (see Theorem 9.1). We choose this specific lattice because it satisfies the conditions of our main Theorem 8.2 and is the unique unimodular even lattice with only one $1$-dimensional cusp. Hence, in this case, the symplectic Eisenstein series corresponds to the full $\textup{Sp}_{2}(\mathbb{Z})$ and therefore a functional equation for it is available.
The structure of the paper is as follows: In Sections 2 and 3, we give the definitions of the main objects of consideration, i.e. modular forms, Fourier-Jacobi forms and the Dirichlet series of interest. In Section 4, we obtain the integral representation of the Dirichlet series, and in Section 5, we rewrite the Eisenstein series of Klingen type in the form of an Epstein zeta function. In Section 6, we define the appropriate theta series for our case, and Section 7 is devoted to discussing the differential operators needed in order to obtain the theta-correspondence. This is done in Section 8. Finally, in Section 9, we obtain a functional equation in the case of the $E_{8}$ lattice.
@@
We denote the space of $m\times n$ matrices with coefficients in a ring $R$ with $\textup{Mat}_{m,n}(R)$. If $n=m$, we often use the notation $\textup{Mat}_{n}(R)$. For any matrix $M\in\textup{Mat}_{n}(R)$, we denote by $\det(M),\textup{ tr}(M)$ the determinant and trace of $M$ respectively. By $\textup{GL}_{n}(R)$, we denote the matrices in $M_{n}(R)$ with non-zero determinant. By $1_{n}$, we denote the $n\times n$ identity matrix. For any vector $v$, we denote by $v^{t}$ its transpose. We also use the bracket notation $A[B]:=\overline{B}^{t}AB$ for suitably sized complex matrices $A,B$. For a complex number $z$, we denote by $e(z):=e^{2\pi iz}$. Finally, let $\zeta(s)$ denote the Riemann zeta function, $\xi(s):=\pi^{-s/2}\Gamma(s/2)\zeta(s)$ denote the completed zeta function and $\Gamma(s):=\int_{0}^{\infty}t^{s-1}e^{-t}\textup{d}t$ denote the Gamma function.
@@ 2. Preliminaries
Let $V$ denote a finite-dimensional vector space over $\mathbb{Q}$. We start with the following Definition.
@@
A $\mathbb{Z}$-lattice $\Lambda$ is a free, finitely generated $\mathbb{Z}$-module, which spans $V$ over $\mathbb{Q}$.
In the following, let $V:=\mathbb{Q}^{n}$ and $L:=\mathbb{Z}^{n}$ with $n\geq 1$. Assume $S$ is an even integral positive definite symmetric matrix of rank $n$. Here, even means $S[x]\in 2\mathbb{Z}$ for all $x\in L$. Let also $Q:=S/2$. We define
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\ S_{1}:=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
of real signatures $(1,n+1)$ and $(2,n+2)$ respectively and denote by $Q_{i}:=S_{i}/2$ for $i=0,1$. Let also $V_{0}:=\mathbb{Q}^{n+2}$ and $V_{1}:=\mathbb{Q}^{n+4}$ and consider the quadratic spaces $(V_{0},\phi_{0}),\ (V_{1},\phi_{1})$, where
$\displaystyle\phi_{i}:V_{i}\times V_{i}$ $\displaystyle\longmapsto\mathbb{Q}$
$\displaystyle(x,y)$ $\displaystyle\longmapsto\frac{1}{2}x^{t}S_{i}y,$
for $i=0,1$. We then have that $\phi:=\phi_{0}\mid_{V\times V}$ is just $(x,y)\longmapsto-x^{t}Sy/2$, and we make the assumption that $L=\mathbb{Z}^{n}$ is a maximal $\mathbb{Z}$-lattice with respect to $\phi$. Here, maximal means that the lattice $L$ is maximal among all $\mathbb{Z}$-lattices in $V$ on which $\phi[x]$ takes values in $\mathbb{Z}$, where we use the notation $\phi[x]:=\phi(x,x)$ for $x\in V$.
From [shimura2004arithmetic, Lemma 6.3], we then obtain that $L_{0}:=\mathbb{Z}^{n+2}$ is a $\mathbb{Z}$-maximal lattice in $V_{0}$.
If now $K\supset\mathbb{Q}$ is a field, we define the corresponding special orthogonal groups of $K$-rational points via
$G^{*}_{K}:=\{g\in\textup{SL}_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},$
$G_{K}:=\{g\in\textup{SL}_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
We view $G^{*}_{K}$ as a subgroup of $G_{K}$ via the embedding $g\longmapsto\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}$.
We also use the notation $\textup{SO}^{\phi_{0}}(V_{0})$ and in general $\textup{SO}^{\phi}(V)$ for any quadratic space $(V,\phi)$.
@@
For a matrix $g\in\textup{Mat}_{n+4}(\mathbb{R})$, we write it as
@@
with $A\in\textup{Mat}_{n+2}(\mathbb{R}),\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $a,b,c,d$ real column vectors. Now the map
(2) $Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$
@@
Let also $\Gamma^{+}(L_{0}):=\Gamma(L_{0})\cap G^{*,0}_{\mathbb{R}}$. If then
@@
where $Z\in\mathcal{H}_{S}$. It is then Koecher’s principle that gives us that $A(r)=0$ unless $r\in L_{0}^{*}\cap\overline{\mathcal{P_{S}}}$ (cf. [eisenstein_thesis, Theorem 1.5.2]). Here, $\overline{\mathcal{P}_{S}}$ denotes the closure of $\mathcal{P_{S}}$. By [eisenstein_thesis, Theorem 1.6.23], and because $L$ is $\mathbb{Z}$-maximal, we have the following definition for cusp forms.
@@
$\langle F,G\rangle:=\int_{\mathcal{Q}}F(Z)\overline{G(Z)}\left(Q_{0}[\textup{Im}Z]\right)^{k}\textup{d}^{*}Z,$
where $\textup{d}^{*}Z=\left(Q_{0}[\textup{Im}Z]\right)^{-(n+2)}\textup{d}Z$ denotes the $G_{\mathbb{R}}^{0}$-invariant volume element on $\mathcal{H}_{S}$. We remind here again that $Q_{0}=S_{0}/2$.